OpenStudy (anonymous):

write an equadratic equation for the variable x having -sqrt5,3sqrt5 as the solution

OpenStudy (anonymous):

\[-\sqrt{5}, 3\sqrt{5}\]

OpenStudy (akshay_budhkar):

if a and b are the solutions of a quadratic equation then the equation is given by (x-a)(x-b)=0 So you can substitute a and b and simplify, can you do that?

OpenStudy (anonymous):

oh i have to use ax62+bx+c=0

OpenStudy (anonymous):

ax^2+bx+c=0

OpenStudy (akshay_budhkar):

Nope, that are different a and b.. whjat i am talking of are different

OpenStudy (akshay_budhkar):

um. you get it?

OpenStudy (anonymous):

start with \[(x-\sqrt{5})(x-3\sqrt{5})=0\] then multiply out. it will be messy

OpenStudy (anonymous):

oops that was wrong. start with \[(x+\sqrt{5})(x-3\sqrt{5})=0\]

OpenStudy (anonymous):

\[x ^{2}-3\sqrt{5}+\sqrt{5x}-3\sqrt{5}\] is tht it?

jhonyy9 (jhonyy9):

from the second term you have missed the x and the last term will be -3*5=-15 so and now i see that again the 3rd term is wrong because x not is inside radical this will be xsqrt5 right,correct - hope that you will can understanding